OpenStudy (anonymous):

What is the domain and range of 1/-3x+2

OpenStudy (abb0t):

denom ≠ 0 therfore solve for x. which is x = 2/3. x must therefore be greater than or equal to 2/3.

OpenStudy (mathstudent55):

Wait a minute there. If the fraction is 1/(-3x + 2), with -3x + 2 as the denominator, then that denominator cannot equal zero. Which value of x makes -3x + 2 = 0? To find that out, solve -3x + 2 = 0 -3x + 2 = 0 -3x = -2 x = 2/3 So the domain is all values of x except x = 2/3

OpenStudy (abb0t):

yep